### v1

# On the critical parameters of branching random walks

Abstract.
Given a discrete spatial structure $X$, we define continuous-time branching processes that model a population breeding and dying on $X$. These processes are usually called branching random walks. They are characterized by breeding rates $k_{xy}$ between sites (ruling the speed at which individuals at $x$ send children to $y$) and by the multiplicative speed parameter $\lambda$. Two critical parameters of interest are the global critical parameter $\lambda_{w}$, related to global survival, and the local critical parameter $\lambda_{s}$, related to survival in finite sets (with $\lambda_{w}\leq\lambda_{s})$. Local modifications of the rates of the process can affect the value of both parameters. Using results on the comparison of the probability of extinction of a single branching random walk on different sets, we are able to extend the comparison to extinction probabilities and critical parameters of couples of branching random walks, whose rates coincide outside a fixed set $A\subseteq X$. We say that two branching random walks are equivalent if their rates coincide everywhere but for a finite subset of $X$. Given an equivalence class of branching random walks, we prove that if one process has $\lambda^{*}_{w}\neq\lambda^{*}_{s}$, then $\lambda_{w}^{*}$ is the maximal possible value of this parameter in the class. We describe the possible situations for the critical parameters in these equivalence classes.
Keywords: branching random walk, branching process, critical parameters, local survival, global survival, pure global survival phase.
AMS subject classification: 60J05, 60J80.
1. Introduction
A branching process, or Galton-Watson process (see [12]), is a process where a particle dies and gives birth to a random number of offspring, according to a given offspring law $\rho$ ($\rho(n)$ being the probability of having exactly $n$ children). Different particles breed independently, according to $\rho$. The process can either go extinct (i.e. no particles are alive from a certain time on) almost surely, or survive (i.e. at least one particle in any generation) with positive probability. The probability of extinction can be computed as a function of $\rho$.
A branching random walk (BRW hereafter) is a process where particles are described by their location $x\in X$, where $X$ is an at most countable set ($X$ is usually interpreted as a spatial variable, but can also be seen as a “type”, see for instance [14]). The particles at site $x\in X$ are replaced by a random number of children, which are placed at various locations on $X$. The breeding law depends on the site where the parent lives and all particles breed independently. This class of processes (in continuous and discrete time) has been studied by many authors (see [1, 11, 13, 15, 16, 17, 18, 19, 20] just to mention a few); a survey on the subject can be found in [7]. Note that, in the case of the branching random walk, there is no upper bound for the number of particles per site. When such an upper bound is fixed, say at most $m$ particles per site, we get the $m$-type contact process. The branching random walk can be obtained as the limiting process as $m$ goes to infinity ([6, 9, 23]).
BRWs can be considered in discrete time, where each particle is alive for one generation and its offspring all live in the next generation, or in continuous time, where generations overlap since particles may breed at different times during their whole lifetime. In particular, one can consider a family of parametrized processes, where we fix the breeding rates between locations and the parameter $\lambda>0$ represent the reproductive speed (the larger $\lambda$, the shorter the time intervals between subsequent breedings, see Section 2 for details).
Since the BRW lives on a spatial structure, its behaviour is in general more complex than the one of a branching process: indeed survival and extinction can studied globally (on the whole space $X$) but also locally (on a single site or a finite set $A$). Starting the process with one particle at a given site $x$, and fixing a set $A\subseteq X$ only one of the following holds for the BRW: (1) it goes almost surely extinct, (2) it survives globally but not locally, (3) it survives globally and locally but with different probabilities (non-strong local survival), (4) it survives globally and locally with equal probability (strong local survival). We stress that no strong local survival means that either there is non-strong local survival or almost sure local extinction.
When considering BRWs in continuous time, a fixed family may exhibit different behaviours for different values of the speed parameter $\lambda$, and the threshold separating global extinction from global survival is called global critical parameter $\lambda_{w}$. Analogously, the threshold between local extinction and local survival is called local critical parameter $\lambda_{s}$. We note that the subscripts $w$ and $s$ stand for ”weak” and ”strong” which are sometimes used as synonyms for ”global” and ”local”, respectively.
The present paper is devoted to the question of what can be said when a BRW is modified locally: this means that the breeding laws are modified only in a fixed set $A\subseteq X$.
For instance we may wonder about what can be said about the modified BRW, if the original BRW has no strong local survival. Theorem 3.2 shows that there is global survival and no strong local survival in $A$ in the original BRW if and only if there is global survival and no strong local survival in $A$ in the modified BRW (regardless of the modifications that have been introduced in $A$). As a corollary we get that if the original BRW dies out locally in $A$ and the modified BRW survives globally, then almost sure global extinction for the original one is equivalent to strong local survival in $A$ for the modified BRW (Corollary 3.4 and Figure 1). Moreover, for a fixed irreducible BRW, if there is global survival and no strong local survival in some $A\subseteq X$ then there is global survival and no strong local survival in all finite $B\subseteq X$.
From these results in discrete time, we are able to prove that, in continuous time, a modification of the BRW in a finite subset $A$, which lowers the global critical parameter (this can be usually achieved by adding a sufficiently fast reproduction rate at some site), implies that the global and local parameters of the modified BRW coincide (Corollary 3.3). This allows us to describe a general method to produce examples such as Example 4.2, where the modified BRW has strong local survival for some values of the parameter below a threshold and above another threshold, but non-strong local survival for intermediate values of the parameter (see Figures 1 and 2). This example was originally described in [8] but appears here with an easier proof and in a more general framework. Moreover, we prove that in general, a continuous-time BRW which is obtained by a local modification of another BRW, lowering its global critical parameter, dies out globally at the global critical parameter (which is not always true, see [5, Example 3]).
Here is the outline of the paper. In Section 2 we recall the definitions of discrete-time and continuous-time BRWs, of the extinction probability vectors and of their asymptotical behaviour (namely, if they go extinct or not as time grows to infinity). Global, local and strong local survival are defined, along with the critical parameters of continuous-time BRWs. Theorem 2.4 and Corollary 2.5 state known results about the approximation of a BRW through a sequence of BRWs, in a form which is useful in the sequel. Proposition 3.6 gives a sufficient condition for the equality $\lambda_{w}=\lambda_{s}$. In Section 3 we first compare the probabilities of extinction of a single BRW on different sets (Theorem 3.1), then the probabilities of extinction of two different BRWs (Theorem 3.2). Afterwards, we proceed to study what happens to the critical parameters of a continuous-time BRW, when it is subjected to a local modification of its rates (Corollaries 3.3, 3.4, 3.5). Section 4 starts from the observation that if a BRW has a pure global survival phase ($\lambda_{w}<\lambda_{s}$, that is, there is a region for the speed parameter $\lambda$ where the process survives globally but eventually leaves all finite sets), then no local modification of the rates can increase the global critical parameter, see Proposition 4.1. This means that the global critical parameter of a BRW with pure global survival phase is maximal, among the same parameters of any BRW obtained through local modification. In Section 4 we analyze the behaviour of the BRW on the homogeneous tree, where we add one reproductive loop at the origin (Example 4.2) and answer several questions in general settings, about the critical parameters, when the rates are modified in a subset $A$.
2. Basic definitions and preliminaries
2.1. Discrete-time Branching Random Walks and their survival/extinction.
Given an at most countable set $X$, we define a discrete-time BRW as a process $\{\eta_{n}\}_{n\in\mathbb{N}}$, where $\eta_{n}(x)$ is the number of particles alive at $x\in X$ at time $n$. The dynamics is described as follows: let $S_{X}:=\{f:X\to\mathbb{N}\colon\sum_{y}f(y)<\infty\}$ and let $\mu=\{\mu_{x}\}_{x\in X}$ be a family of probability measures on the (countable) measurable space $(S_{X},2^{S_{X}})$. A particle of generation $n$ at site $x\in X$ lives one unit of time; after that, a function $f\in S_{X}$ is chosen at random according to the law $\mu_{x}$. This function describes the number of children and their positions, that is, the original particle is replaced by $f(y)$ particles at $y$, for all $y\in X$. The choice of $f$ is independent for all breeding particles. The BRW is denoted by $(X,\mu)$.
Some results rely on the first-moment matrix $M=(m_{xy})_{x,y\in X}$, where each entry $m_{xy}:=\sum_{f\in S_{X}}f(y)\mu_{x}(f)$ represents the expected number of children that a particle living at $x$ sends to $y$ (briefly, the expected number of particles from $x$ to $y$). For the sake of simplicity, we require that $\sup_{x\in X}\sum_{y\in X}m_{xy}<+\infty$.
[…]
We consider initial configurations with only one particle placed at a fixed site $x$ and we denote by $\mathbb{P}^{x}$ the law of the corresponding process. The evolution of the process with more than one initial particle can be obtained by superimposition.
In the following, wpp is shorthand for “with positive probability” (although, when talking about survival, “wpp” will be usually tacitly understood). In order to avoid trivial situations where particles have one offspring almost surely, we assume henceforth the following.
Assumption 2.1.
[…]
We define the extinction probabilities of the process and the possible behaviours of the BRW.
Definition 2.2.
Given a BRW $(X,\mu)$, $x\in X$ and $A\subseteq X$, the probability of extinction in $A$, starting from one particle at $x$, is defined as
${\mathbf{q}}(x,A):=1-\mathbb{P}^{x}(\limsup_{n\to\infty}\sum_{y\in A}\eta_{n}(y)>0).$
We denote by ${\mathbf{q}}(A)$ the extinction probability vector, whose $x$-entry is ${\mathbf{q}}(x,A)$.
Note that ${\mathbf{q}}(x,A)$ depend on $\mu$. When we need to stress this dependence, we write ${\mathbf{q}}^{\mu}(x,A)$. Extinction probabilities have been the object of intense study during the last decades. We refer the reader, for instance, to [2, 7, 23].
Definition 2.3.
Let $(X,\mu)$ be a BRW starting from one particle at $x\in X$ and let $A\subseteq X$. We say that
  1. (1)
    the process survives locally wpp in $A$ if ${\mathbf{q}}(x,A)<1$;
  2. (2)
    the process survives globally wpp if ${\mathbf{q}}(x,X)<1$;
  3. (3)
    there is strong local survival wpp in $A$ if ${\mathbf{q}}(x,A)={\mathbf{q}}(x,X)<1$ and non-strong local survival wpp in $A$ if ${\mathbf{q}}(x,X)<{\mathbf{q}}(x,A)<1$;
  4. (4)
    the process is in a pure global survival phase starting from $x$ if ${\mathbf{q}}(x,X)<{\mathbf{q}}(x,x)=1$, (where we write ${\mathbf{q}}(x,y)$ instead of ${\mathbf{q}}(x,\{y\})$ for all $x,y\in X$).
When there is no survival wpp, we say that there is extinction and the fact that extinction occurs almost surely will be tacitly understood.
When there is strong local survival, it means that for almost all realizations the process either survives locally (hence globally) or it goes globally extinct. More precisely, there is strong survival at $y$ starting from $x$ if and only if the probability of local survival at $y$ starting from $x$ conditioned on global survival starting from $x$ is $1$.
We want to stress that, for a BRW starting from $x\in X$, ${\mathbf{q}}(x,X)={\mathbf{q}}(x,A)$ if and only if global survival is equivalent to strong local survival at $A$. On the other hand ${\mathbf{q}}(x,X)<{\mathbf{q}}(x,A)$ if and only if there is global survival and no strong local survival at $A$ (that is, either local extinction at $A$ or non-strong local survival at $A$). Recall that no strong local survival in $A$ means that either there is non-strong local survival in $A$ or there is local extinction in $A$.
2.2. Continuous-time Branching Random Walks
Given a graph $(X,E)$ and a nonnegative matrix $K=(k_{xy})_{(x,y)\in E}$, one can define a family of continuous-time Branching Random Walks $\{\eta_{t}\}_{t\geq 0}$, where $\eta_{t}(x)$ represents the number of particles alive at time $t$ at site $x$, for any $x\in X$. The family is indexed by the reproductive speed parameter $\lambda>0$.
Each particle has an exponentially distributed lifetime with parameter 1. During its lifetime each particle alive at $x$ breeds into $y$ according to the arrival times of its own Poisson process with parameter $\lambda k_{xy}$ (representing the reproduction rate). We denote by $(X,K)$ this family of continuous-time BRWs (depending on $\lambda>0$). It is not difficult to see that the introduction of a nonconstant death rate $\{d(x)\}_{x\in X}$ does not represent a significant generalization. Indeed, one can study a new BRW with death rate 1 and reproduction rates $\{\lambda k_{xy}/d(x)\}_{x,y\in X}$; the two processes have the same behaviours in terms of survival and extinction ([8, Remark 2.1]).
To each continuous-time BRWs we associate a discrete-time counterpart, namely $(X,\mu)$ where we simply take into account the number and the positions of all the offsprings born before the death of the parent. Clearly, from the knowledge of the discrete-time counterpart of $(X,K)$, we cannot retrieve how many particles are alive at time $t$ (and where), but the information about survival and extinction is intact. In particular, by definition, a continuous-time BRW has some property if and only if its discrete-time counterpart has it.
Given $x\in X$, two critical parameters are associated to the continuous-time BRW: the global survival critical parameter $\lambda_{w}(x)$ and the local survival critical parameter $\lambda_{s}(x)$ (or, briefly, the global and local critical parameters, respectively) defined as
$\begin{split}\lambda_{w}(x)&:=\inf\Big{\{}\lambda>0\colon\,\mathbb{P}^{x}\Big{(}\sum_{w\in X}\eta_{t}(w)>0,\forall t\Big{)}>0\Big{\}},\\ \lambda_{s}(x)&:=\inf\{\lambda>0\colon\,\mathbb{P}^{x}\big{(}\limsup_{t\to\infty}\eta_{t}(x)>0\big{)}>0\}.\end{split}$
These values depend only on the irreducible class of $x$; in particular they are constant if the BRW is irreducible.
Whenever the interval $(\lambda_{w}(x),\lambda_{s}(x))$ is non-empty we say that there exists a pure global survival phase starting from $x$. No reasonable definition of a strong local survival critical parameter is possible (see [8]).
It is possible to identify $\lambda_{s}(x)$ in terms of the first moment matrix $M$. Indeed, consider the $n$th power matrix $M^{n}$ with entries $m^{(n)}_{xy}$: $\lambda_{s}(x)=1/\limsup_{n\in\mathbb{N}}\sqrt[n]{m^{(n)}_{xx}}$ (see [21]). One can also write the local critical parameter in terms of a generating function. Let
$\begin{split}\varphi^{(n)}_{xy}&:=\sum_{x_{1},\ldots,x_{n-1}\in X\setminus\{y\}}m_{xx_{1}}m_{x_{1}x_{2}}\cdots m_{x_{n-1}y}\\ \Phi(x,y|\lambda)&:=\sum_{n=1}^{\infty}\varphi_{xy}^{(n)}\lambda^{n}.\end{split}$
The coefficients $\varphi^{(n)}_{xy}$ play in the BRW theory the same role played by the first-return probabilities in random walk theory (see [22, Section 1.C]). Roughly speaking, $\varphi^{(n)}_{xy}$ is the expected number of particles alive at $y$ at time $n$, when the initial state is just one particle at $x$ and the process behaves like a BRW except that every particle reaching $y$ at any time $i<n$ is immediately killed (before breeding). The following characterization holds (see [4, 5]):
$\lambda_{s}(x)=\max\{\lambda\in{\mathbb{R}}:\Phi(x,x|\lambda)\leq 1\}.$
The characterization of $\lambda_{w}(x)$ is less explicit: if we denote by $l_{*}^{\infty}(X)$ the usual subset of nonnegative, bounded functions on $X$, we have (see [5])
$\lambda_{w}(x)=\inf\{\lambda\in\mathbb{R}:\exists v\in l^{\infty}_{+}(X),\,v(x)>0\textrm{ and }n\in\mathbb{N},\,\lambda^{n}K^{n}v\geq v\}$
where $K^{n}v(x):=\sum_{y\in X}k^{(n)}_{xy}v(y)$ ($k^{(n)}_{xy}$ are the entries of the matrix $K^{n}$, the $n$th power of the matrix $N$). More precisely, there is global survival for a fixed $\lambda>0$ if and only if $\frac{\lambda Kv(x)}{1+\lambda Kv(x)}\geq v(y)$ for all $x\in X$.
2.3. Spatial Approximation
Given a sequence of BRWs $\{(X_{n},\mu_{n})\}_{n\in\mathbb{N}}$ such that $\liminf_{n\to\infty}X_{n}=X$, we define $m(n)_{xy}:=\sum_{f\in S_{X_{n}}}f(y)\mu_{n,x}(f)$, where $x,y\in X_{n}$, and the corresponding sequence of first-moment matrices $\{M_{n}\}_{n\in\mathbb{N}}$. Clearly, for all $x,y\in X$, $m(n)_{xy}$ is eventually well defined as $n\to+\infty$. Indeed, $\liminf_{n\to\infty}X_{n}=X$, implies that given $A\subseteq X$ finite, $A\subseteq X_{n}$ for every sufficiently large $n$. We are wondering when a BRW can be approximated by a sequence of BRWs: more precisely, we are focusing on the local behaviour. Note that in the following results we are not assuming that the BRW is irreducible (cfr. [7, Section 5.1], [23, Theorem 5.2] and [5, Theorem 3.1]).
Theorem 2.4.
Let us fix a vertex $x_{0}\in X$. Suppose that $\liminf_{n\to\infty}X_{n}=X$ and $m(n)_{xy}\leq m_{xy}$ for all $x,y\in X_{n}$ and all $n\in\mathbb{N}$. Assume that, for all $x,y\in X$, $m(n)_{xy}\to m_{xy}$ as $n\to\infty$.
  1. (1)
    if $(X,\mu)$ dies out locally (resp. globally) a.s. starting from $x_{0}$ then $(X_{n},\mu_{n})$ dies out locally (resp. globally) a.s starting from $x_{0}$ for all $n\in\mathbb{N}$;
  2. (2)
    if $(X,\mu)$ survives locally starting from $x_{0}$ then $(X_{n},\mu_{n})$ survives locally starting from $x_{0}$ eventually as $n\to\infty$.
Theorem 2.4 yields an analogous result for BRWs in continuous time.
Corollary 2.5.
Let $(X,K)$ be a continuous-time BRW and let us consider a sequence of continuous-time BRWs $\{(X_{n},{K_{n}})\}_{n\in\mathbb{N}}$ such that $\liminf_{n\to\infty}X_{n}=X$. Let us suppose that $k_{xy}(n)\leq k_{xy}$ for all $n\in\mathbb{N}$, $x,y\in X_{n}$ and $k_{xy}(n)\to k_{xy}$ as $n\to\infty$ for all $x,y\in X$. Then $\lambda_{s}^{(X_{n},{K_{n}})}(x_{0})\geq\lambda_{s}^{(X,K)}(x_{0})$ for every $n\in\mathbb{N}$ and $x_{0}\in X_{n}$. Moreover, given $x_{0}\in X$, we have $x_{0}\in X_{n}$ eventually as $n\to+\infty$ and $\lambda_{s}^{(X_{n},{K_{n}})}(x_{0})\to\lambda_{s}^{(X,K)}(x_{0})$ as $n\to\infty$.
Among all possible choices of the sequence $\{(X_{n},\mu_{n})\}_{n\in\mathbb{N}}$ there is one which is induced by $(X,\mu)$ on the subsets $\{X_{n}\}_{n\in\mathbb{N}}$; more precisely, one can take $\mu_{n}(g):=\sum_{f\in S_{X}:f|_{X_{n}}=g}\mu_{x}(f)$ for all $x\in X_{n}$ and $g\in S_{X_{n}}$. Roughly speaking, this choice means that all reproductions outside $X_{n}$ are suppressed. In this case it is simply $m(n)_{xy}=m_{xy}$ for all $x,y\in X_{n}$.
3. Strong local survival and local modifications
We recall here [2, Theorem 4.1], which is a fundamental tool for comparing extinction probability vectors. In the case of global survival (choose $B=X$), it gives equivalent conditions for strong local survival in terms of extinction probabilities. We denote by ${\mathbf{q}}_{0}({x},A)$ the probability that the process, starting from one particle at $x$, has no progeny ever in the set $A$.
Theorem 3.1.
For any BRW and $A,B\subseteq X$, the following statements are equivalent:
  1. (1)
    there exists $x\in{X}$ such that ${\mathbf{q}}(x,B)<{\mathbf{q}}({x},A)$;
  2. (2)
    there exists $x\in{X}$ such that ${\mathbf{q}}({x},B\setminus A)<{\mathbf{q}}({x},A)$;
  3. (3)
    there exists $x\in{X}$ such that ${\mathbf{q}}({x},B)<{\mathbf{q}}_{0}({x},A)$;
  4. (4)
    there exists $x\in{X}$ such that, starting from $x$ there is a positive chance of survival in $B$ without ever visiting $A$;
  5. (5)
    there exists $x\in{X}$ such that, starting from $x$ there is a positive chance of survival in $B$ and extinction in $A$;
  6. (6)
    $\inf_{x\in{X}\colon{\mathbf{q}}(x,B)<1}\frac{1-{\mathbf{q}}(x,A)}{1-{\mathbf{q}}(x,B)}=0.$
From Theorem 3.1, which is stated for a single BRW, we derive Theorem 3.2 and Corollary 3.3 which give us information about the behaviour of a BRW after some modifications.
To this aim, consider two BRWs $(X,\mu)$ and $(X,\nu)$ and denote by ${\mathbf{q}}^{\mu}$ and ${\mathbf{q}}^{\nu}$ their respective extinction probability vectors. If $\mu_{x}=\nu_{x}$ for all $x\not\in A$, for some set $A\subseteq X$, we derive properties of ${\mathbf{q}}^{\nu}$ from analogous properties of ${\mathbf{q}}^{\mu}$.
Theorem 3.2.
Let $(X,\mu)$ and $(X,\nu)$ be two BRWs. Suppose that $A\subseteq X$ is a nonempty set such that $\mu_{x}=\nu_{x}$ for all $x\not\in A$. Then we have that ${\mathbf{q}}_{0}^{\mu}(x,A)={\mathbf{q}}_{0}^{\nu}(x,A)$ for all $x\in X$ and, for all $B\subseteq X$,
${\mathbf{q}}^{\mu}(A)\leq{\mathbf{q}}^{\mu}(B)\Longleftrightarrow{\mathbf{q}}^{\nu}(A)\leq{\mathbf{q}}^{\nu}(B).$
If, in addition, $\mu_{x}=\nu_{x}$ for all $x\not\in B$, then
${\mathbf{q}}^{\mu}(A)={\mathbf{q}}^{\mu}(B)\Longleftrightarrow{\mathbf{q}}^{\nu}(A)={\mathbf{q}}^{\nu}(B).$
Proof.
We note that $(X,\mu)$ and $(X,\nu)$ have the same behaviour until they first hit $A$, hence ${\mathbf{q}}_{0}^{\mu}(x,A)={\mathbf{q}}_{0}^{\nu}(x,A)$ for all $x\not\in A$. If $x\in A$, then clearly ${\mathbf{q}}_{0}^{\mu}(x,A)=0={\mathbf{q}}_{0}^{\nu}(x,A)$.
Suppose now that ${\mathbf{q}}^{\mu}(A)\nleq{\mathbf{q}}^{\mu}(B)$. Hence, according to Theorem 3.1, there exists $x\in X\setminus A$ such that there is a positive probability of survival in $B$ starting from $x$ without ever visiting $A$. Since the two processes have the same behaviour until they first hit $A$, the same holds for $(X,\nu)$ and this implies that ${\mathbf{q}}^{\nu}(x,A)>{\mathbf{q}}^{\nu}(x,B)$; thus ${\mathbf{q}}^{\nu}(A)\nleq{\mathbf{q}}^{\nu}(B)$. The reversed implication follows by exchanging the role of $\mu$ and $\nu$.
The final double implication, ${\mathbf{q}}^{\mu}(A)={\mathbf{q}}^{\mu}(B)\Longleftrightarrow{\mathbf{q}}^{\nu}(A)={\mathbf{q}}^{\nu}(B),$ follows from the previous one by exchanging the role of $A$ and $B$. ∎
The previous result extends [10, Theorem 4.1] by considering a generic subset $B$ instead of $X$. In particular, if $A_{0}:=\{x\in X\colon\mu_{x}\neq\nu_{x}\}$ then the partial order of the sets $\{\mathbf{q}^{i}(A)\colon A_{0}\subseteq A\subseteq X\}$ ($i=\mu,\,\nu$) is preserved by the (well-defined) map $\mathbf{q}^{\mu}(A)\mapsto\mathbf{q}^{\nu}(A)$.
Theorem 3.2, applied to continuous-time BRWs, leads to the following corollary which describes how local modifications of the BRW affect the critical parameters (see [10, Corollary 4.4]).
Corollary 3.3.
Let $(X,K)$ and $(X,K^{*})$ two irreducible continuous-time BRWs such that $k_{xy}=k^{*}_{xy}$ for all $x\in X\setminus A$ where $A$ is a nonempty, finite set. Then the following are equivalent:
  1. (1)
    $\lambda^{*}_{w}<\lambda_{w}$;
  2. (2)
    $\lambda^{*}_{s}<\lambda_{w}$;
  3. (3)
    $\lambda^{*}_{w}=\lambda^{*}_{s}<\lambda_{w}$.
In particular if $\min(\lambda_{s},\lambda_{s}^{*})\geq\max(\lambda_{w},\lambda_{w}^{*})$ (for instance if $\lambda_{s}=\lambda_{s}^{*}$) then $\lambda_{w}=\lambda_{w}^{*}$.
Proof.
Observe that the discrete-time counterparts of these continuous-time BRWs satisfy the hypotheses of Theorem 3.2, namely, their offspring distribution are the same outside $A$.
Clearly 2.1.2$\Longrightarrow$ 1.2 . ⟹ 1 . and 3.2.3$\Longrightarrow$ 2.3 . ⟹ 2 .. We just need to prove that 1.3.1$\Longrightarrow$ 3.1 . ⟹ 3 .; more precisely, we prove that $\lambda^{*}_{w}<\lambda_{w}\Longrightarrow\lambda^{*}_{w}=\lambda^{*}_{s}$. Take $\lambda\in(\lambda_{w}^{*},\lambda_{w})$; the $\lambda$-$(X,K^{*})$ BRW survives globally, hence ${\mathbf{q}}^{*}(X)<\mathbf{1}$. On the other hand, $\mathbf{1}={\mathbf{q}}(X)=\mathbf{q}(A)$ whence, according to Theorem 3.2(1), ${\mathbf{q}}^{*}(X)=\mathbf{q}^{*}(A)$ which implies $\mathbf{q}^{*}(A)<\mathbf{1}$. If the $\lambda$-$(X,K^{*})$ BRW survives locally in the finite set $A$ it means that it survives locally at a vertex $x\in A$ ($\Longleftrightarrow$ at every vertex, since the process is irreducible). This implies $\lambda\geq\lambda_{s}^{*}$; thus $\lambda_{s}^{*}\leq\lambda_{w}^{*}$ whence $\lambda_{s}^{*}=\lambda_{w}^{*}$. ∎
According to Corollary 3.3, when $(X,K)$ and $(X,K^{*})$ are two irreducible continuous-time BRWs such that $k_{xy}=k^{*}_{xy}$ for all $x\in X\setminus A$ where $A$ is a nonempty, finite set, then the only alternatives are:
  1. (1)
    $\lambda_{w}=\lambda_{w}^{*}\leq\min(\lambda_{s},\lambda_{s}^{*})$,
  2. (2)
    $\lambda_{w}=\lambda_{s}<\lambda_{w}^{*}\leq\lambda_{s}^{*}$,
  3. (3)
    $\lambda_{w}^{*}=\lambda_{s}^{*}<\lambda_{w}\leq\lambda_{s}$.
We are thus able to describe the behaviour of the modified BRW, when the original BRW has a pure weak phase and the modified has no weak phase. The phase diagram can be seen in Figure 1. Note that, while in general a BRW may survive or die out globally at its global critical parameter, here we know the behaviour at $\lambda=\lambda^{*}_{w}$ (the process dies out globally). Moreover, for $\lambda>\lambda_{s}$ there is local survival which is strong or non-strong depending on the behaviour of $(X,K)$. These results are summarized in Corollary 3.4.
[…]
Figure 1. The phase diagram of the BRW $(X,K^{*})$ in finite $B\neq\emptyset$.
Corollary 3.4.
Let $(X,K)$ and $(X,K^{*})$ two irreducible continuous-time BRWs such that $k_{xy}=k^{*}_{xy}$ for all $x\in X\setminus A$ where $A$ is a nonempty, finite set. Suppose that $\lambda^{*}_{w}<\lambda_{w}<\lambda_{s}$, and let $B\subseteq X$ be nonempty. Then, for the BRW $(X,K^{*})$
  1. (1)
    if $\lambda\leq\lambda_{w}^{*}$ there is a.s. extinction in $B$;
  2. (2)
    if $\lambda\in(\lambda_{w}^{*},\lambda_{w})$ there is strong survival in $B$;
  3. (3)
    if $\lambda=\lambda_{w}$ and $(X,K)$ dies out globally, then there is strong survival in $B$; conversely, if $(X,K)$ survives globally and $B$ is finite; there is non-strong survival in $B$;
  4. (4)
    if $\lambda\in(\lambda_{w},\lambda_{s}]$ and $B$ is finite then there is non-strong survival in $B$;
  5. (5)
    if $\lambda>\lambda_{s}$ and $B$ is finite then survival is strong (resp. non-strong) in $B$ if and only if the same holds for $(X,K)$.
Proof.
Note that in the discrete-time counterpart of a continuous-time BRW every particle at every vertex has a positive probability of dying without breeding; hence by [10, Remark 3.1] strong local survival is a common property of all starting vertices. Moreover, if a BRW is irreducible, $A,B\subset X$ are nonempty, finite subsets and $C\subset X$ is nonempty, then $\mathbf{q}(A)=\mathbf{q}(B)\geq\mathbf{q}(C)$. The first equality holds since local survival does not depends on the vertex and survival in a finite (nonempty) set is equivalent to local survival to a vertex in the set. The second inequality follows from the fact that any nonempty set contains a finite nonempty set.
We consider the following disjoint intervals for $\lambda$.
  1. (1)
    Suppose that $\lambda<\lambda_{w}^{*}$; by definition there is global, hence local, extinction. If $\lambda=\lambda_{w}^{*}$ then, according to [5, Theorem 4.7] (see also [4, Theorem 3.5 and Section 4.2]), since $\lambda=\lambda_{w}^{*}=\lambda_{s}^{*}$ then the $\lambda$-$(X,K^{*})$ BRW dies out locally (at any finite set $C$), hence $\mathbf{q}^{*}(C)=\mathbf{1}$ (clearly, being $\lambda<\lambda_{w}$, $\mathbf{q}(B)={\mathbf{q}}(X)=\mathbf{1}$ for all $B\subseteq X$), using Theorem 3.2(1),
    $\mathbf{q}(X)=\mathbf{q}(A)\leq{\mathbf{q}}(B)\Longrightarrow{\mathbf{q}}^{*}(X)=\mathbf{q}^{*}(A)=\mathbf{1}\leq\mathbf{q}^{*}(B).$
    Thus, we have $\mathbf{q}^{*}(B)=\mathbf{1}$.
  2. (2)
    $\lambda\in(\lambda_{w}^{*},\lambda_{w})$. By definition, since $\lambda_{w}^{*}=\lambda_{s}^{*}$, there is global and local survival for the $\lambda$-$(X,K^{*})$ BRW. This implies that ${\mathbf{q}}^{*}(X)\leq\mathbf{q}^{*}(B)<\mathbf{1}$ for every (finite or infinite) set $B$. On the other hand, there is global and local extinction for the $\lambda$-$(X,K)$ BRW which implies ${\mathbf{q}}(X)=\mathbf{q}(B)=\mathbf{1}$. Again, according to Theorem 3.2(1), ${\mathbf{q}}^{*}(X)=\mathbf{q}^{*}(B)$ for every set $B$.
  3. (3)
    Clearly, since $\lambda=\lambda_{w}\leq\lambda_{s}$, we have $\mathbf{q}(B)=\mathbf{1}$ for all finite subsets $B$. Hence
    $\lambda-(X,K)\textrm{ survives globally}\Longleftrightarrow{\mathbf{q}}(X)<\mathbf{q}(A),$
    that is, according to Theorem 3.2(1), if and only if ${\mathbf{q}}^{*}(X)<\mathbf{q}^{*}(A)$. This, again, implies ${\mathbf{q}}^{*}(X)<\mathbf{q}^{*}(B)$ for every nonempty finite subset $B$. If, on the other hand, $\lambda-(X,K)$ dies out globally, then ${\mathbf{q}}^{*}(X)=\mathbf{q}^{*}(A)$ and ${\mathbf{q}}^{*}(X)=\mathbf{q}^{*}(B)$ for every nonempty subset $B$.
  4. (4)
    $\lambda\in(\lambda_{w},\lambda_{s}]$ (we suppose that the interval is nonempty, otherwise there is nothing to prove). Here we have ${\mathbf{q}}(X)<\mathbf{1}=\mathbf{q}(B)$ for every finite subset $B$. Theorem 3.2 yields ${\mathbf{q}}^{*}(X)<\mathbf{q}^{*}(A)<\mathbf{1}$ and, by irreducibility, ${\mathbf{q}}^{*}(X)<\mathbf{q}^{*}(B)<\mathbf{1}$ for every finite, nonempty subset $B$.
  5. (5)
    $\lambda>\lambda_{s}$. Now, $\mathbf{q}(B)<\mathbf{1}$ and $\mathbf{q}^{*}(B)<\mathbf{1}$ for every nonempty $B\subset X$. Again, by Theorem 3.2(1), we have
    ${\mathbf{q}}(A)={\mathbf{q}}(X)\Longleftrightarrow{\mathbf{q}}^{*}(A)={\mathbf{q}}^{*}(X).$
    If $B$ is finite then Theorem 3.2(2) yields the conclusion.
In particular, if we know that $(X,K)$ exhibits non-strong local survival in a set $A$, for a fixed value of $\lambda$, we can deduce the behavior of a BRW, which coincides with $(X,K)$ outside $A$.
Corollary 3.5.
Let $(X,K)$ and $(X,K^{*})$ two irreducible continuous-time BRWs such that $k_{xy}=k^{*}_{xy}$ for all $x\in X\setminus A$ where $A$ is a nonempty (not necessarily finite) set. If there exists $\lambda_{0}$ such that $(X,K)$ survives non-strong locally in $A$, then
  1. (1)
    $(X,K^{*})$ survives globally when $\lambda=\lambda_{0}$;
  2. (2)
    if $A$ is finite and $\lambda^{*}_{s}\geq\lambda_{0}$, then there is pure global survival for $(X,K^{*})$ when $\lambda=\lambda_{0}$.
Proof.
If $\lambda=\lambda_{0}$, $\mathbf{q}(A)>\mathbf{q}(X)$ whence, according to Theorem 3.2, $\mathbf{q}^{*}(A)>\mathbf{q}^{*}(X)$; this implies (1) that $\mathbf{q}^{*}(X)<\mathbf{1}$ (hence, there is global survival) and (2) if $\lambda^{*}_{s}\geq\lambda_{0}$ and $A$ is finite then $\mathbf{1}=\mathbf{q}(\{x\})=\mathbf{q}(A)>\mathbf{q}(X)$, where $x$ is a generic point (hence there is pure global survival phase). ∎
A simple class of irreducible BRWs where $\lambda_{w}=\lambda_{s}$ is described by the following result. Here we consider the Alexandroff one-point compactification of $X$ endowed by the discrete topology: the neighbourhoods of $\infty$ are $\{X\setminus A\colon A\subset X,\textrm{ finite}\}$. The $\limsup_{x\to\infty}$ is defined accordingly.
Proposition 3.6.
Let $(X,K)$ be an irreducible BRW such that $\lambda_{s}\limsup_{x\to+\infty}\sum_{y}k_{xy}\leq 1$; then $\lambda_{s}=\lambda_{w}$.
Proof.
By hypotheses there exists $\alpha\geq 0$ such that (1) $\lambda_{s}\alpha\leq 1$ and (2) for every $\varepsilon>0$ there exists a finite $A=A_{\varepsilon}\subseteq X$ such that $\sup_{x\in X\setminus A}\sum_{y}k_{xy}<\alpha+\varepsilon$. Suppose, by contradiction, that $\lambda_{w}<\lambda_{s}$ and let $\lambda\in(\lambda_{w},\lambda_{s}]$; for every $\varepsilon>0$, since there is a.s. extinction in every finite set, by Theorem 3.1 there is a positive probability of global survival starting from some $x\in X\setminus A_{\varepsilon}$ without visiting $A_{\varepsilon}$. Since the total number of particles generated outside $A_{\varepsilon}$ is dominated by a branching process with reproduction rate $\lambda(\alpha+\varepsilon)$, we have that $\lambda(\alpha+\varepsilon)>1$. Since this is true for every $\varepsilon>0$ whe have proven that $\lambda\in(\lambda_{w},\lambda_{s}]$ implies $\lambda\alpha\geq 1$. We have two cases:
  1. (1)
    if $\alpha=0$ then $\lambda\alpha<1$ therefore the set $(\lambda_{w},\lambda_{s}]$ is empty, that is, $\lambda_{s}=\lambda_{w}$;
  2. (2)
    if $\alpha>0$ then $\lambda_{s}\alpha\leq 1$ and $\lambda\in(\lambda_{w},\lambda_{s}]$ imply $\lambda_{w}\alpha\geq 1$ which, in turn, implies $\lambda_{w}\geq\lambda_{s}$; thus $\lambda_{w}=\lambda_{s}$.
4. Maximality of the pure global survival phase
Let us consider the set of irreducible continuous time BRWs on $X$, namely $\mathcal{X}:=\{(X,K)\colon K=(k_{xy})_{x,y\in X}\text{ is irreducible}\}$. Define the relation $\mathcal{R}$ on $\mathcal{X}$ by $(X,K)\,\mathcal{R}\,(X,K^{*})$ if and only if there exists a finite $A\subset X$ such that $k_{xy}=k^{*}_{xy}$ for all $x\in X\setminus A$ and all $y\in X$. An equivalent definition is $(X,K)\,\mathcal{R}\,(X,K^{*})$ if and only if the set $\Delta_{K,K^{*}}:=\{x\in X\colon\exists y\in X,\,k_{xy}\neq k^{*}_{xy}\}$ is finite.
Clearly $\Delta_{K,K}=\emptyset$, whence $\mathcal{R}$ is reflexive; moreover $\Delta_{K,K^{*}}=\Delta_{K^{*},K}$ therefore $\mathcal{R}$ is symmetric. Finally for all $K,K^{*},\widetilde{K}$ we have $\Delta_{K,K^{*}}\subseteq\Delta_{K,\widetilde{K}}\cup\Delta_{\widetilde{K},K^{*}}$ which implies that $\mathcal{R}$ is transitive. The relation $\mathcal{R}$ is an equivalence relation. To avoid a cumbersome notation, we denote by $[K]$ the equivalence class of $(X,K)$ and by $\mathcal{X}/_{\mathcal{R}}$ the quotient set. Henceforth, given an irreducible BRW $(X,K)$ we denote by $\lambda_{w}^{K}$ and $\lambda_{s}^{K}$ its global and local critical values respectively.
Proposition 4.1.
Let $(X,K)\in\mathcal{X}$ such that $\lambda_{w}^{K}<\lambda_{s}^{K}$. Then for all $K^{*}\in[K]$ we have $\lambda_{w}^{K^{*}}\leq\lambda_{w}^{K}$. Moreover, for all $K^{*}\in[K]$ such that $\lambda_{w}^{K^{*}}<\lambda_{s}^{K^{*}}$, we have $\lambda_{w}^{K^{*}}=\lambda_{w}^{K}$.
Proof.
By contradiction, if $\lambda_{w}^{K^{*}}>\lambda_{w}^{K}$ then, according to Corollary 3.3, we have $\lambda_{s}^{K}=\lambda_{w}^{K}$, which contradicts the hypothesis. From the previous part, it follows easily that if $\lambda_{w}^{K}<\lambda_{s}^{K}$ and $\lambda_{w}^{K^{*}}<\lambda_{s}^{K^{*}}$ then $\lambda_{w}^{K^{*}}=\lambda_{w}^{K}$. ∎
As an application of Proposition 4.1, we discuss the behaviour of the critical parameters of the BRW on the $d$-dimensional regular tree $\mathbb{T}_{d}$, when we add the possibility of reproduction from the origin of the tree to itself.
Example 4.2.
Consider the BRW on $X=\mathbb{T}_{d}$, where $k_{xy}=1$ for all couples of neighbours $(x,y)$. For this BRW, $\lambda_{s}=1/2\sqrt{d-1}$ and $\lambda_{w}=1/d$ and the process exhibits a pure global survival phase. We modify this BRW by adding a loop at the origin $o$, and choosing $k_{oo}>0$. We denote by $\lambda_{w}^{*}$ and $\lambda_{s}^{*}$ the global and local critical parameters, respectively, of this modified process. It is well known that for small values of $k_{oo}$, $\lambda_{w}^{*}<\lambda_{s}^{*}$, while for $k_{oo}$ sufficiently large, the critical parameters coincide (see for instance [8, Example 4.2]). We are now able to prove that for $k_{oo}\in[0,(d-2)/\sqrt{d-1}]$, the two parameters coincide with the parameters of the original process ($k_{oo}=0$); for $k_{oo}\in((d-2)/\sqrt{d-1},d(d-2)/(d-1))$ we have $\lambda_{w}^{*}=\lambda_{w}$, while $\lambda_{s}^{*}\in(\lambda_{w},\lambda_{s})$ and is monotonically decreasing. When $k_{oo}\geq d(d-2)/(d-1))$ the two parameters coincide and they tend to 0 as $k_{oo}$ goes to infinity. The plot of the two functions $\lambda_{w}^{*}=\lambda_{w}^{*}(k_{oo})$ and $\lambda_{s}^{*}=\lambda_{s}^{*}(k_{oo})$ can be seen in Figure 2 (note that the axis do not share the same scale).
In order to compute the function $\lambda_{s}^{*}=\lambda_{s}^{*}(k_{oo})$, recall that the local critical parameter equals $\max\{\lambda>0\colon\Phi(o,o|\lambda)\leq 1\}$ (see Section 2 or [5]), where
$\Phi(o,o|\lambda)=d\frac{1-\sqrt{1-4(d-1)\lambda^{2}}}{2(d-1)}+{k_{oo}}\lambda$
The computation of the generating function $\Phi$ in this case is simple: there is only one path from $o$ to $o$ of length 1 (the loop) whence $\varphi^{(1)}_{oo}=k_{oo}\lambda$; on the other hand, the loop is not involved in any 1st-return path of length $n\geq 1$ therefore the coefficient $\varphi^{(n)}_{oo}$ for $n\geq 1$ is the same as in the homogeneous tree (that is, when $k_{oo}=0$). Hence, $\Phi$ is the sum of $k_{oo}\lambda$ and the analogous generating function of the homogeneous tree. It is an exercise to obtain that, for $k_{oo}>(d-2)/\sqrt{d-1}$,
$\lambda_{s}^{*}=\frac{(d-2)k_{oo}+d\sqrt{k_{oo}^{2}+4}}{2(d^{2}+(d-1)k_{oo}^{2})}.$
Note that this provides an example where there are several (more precisely, infinitely many) local modifications of a BRW, all with $\lambda_{w}^{*}<\lambda_{s}^{*}$, and by Proposition 4.1 they all share the same global critical parameter. Moreover, one of these modifications (the one with $k_{oo}=d(d-2)/(d-1)$), has no pure global survival phase but still the same global critical parameter.
$k_{oo}$$\lambda_{s}=1/2\sqrt{d-1}$$\lambda_{w}=1/d$$(d-2)/\sqrt{d-1}$$d(d-2)/(d-1)$$y=\lambda_{s}^{*}$$y=\lambda_{w}^{*}$
Figure 2. The critical parameters of the modified BRW on $\mathbb{T}_{d}$
In the following sections we answer some natural questions about these equivalence classes. Let us summarize the results. A class might have BRWs with pure global survival phase or not (see Section 4.1) but every class necessarily has an uncountable number of BRWs with no pure global survival phase (see Section 4.3). If a class has a BRW with pure global survival phase then the supremum of $\lambda_{w}$ is finite and it is attained; in this case there might also be BRWs with no pure global survival phase in the same class with the same maximum $\lambda_{w}$ (see Section 4.4). If a class has only BRWs with no pure global survival phase then the supremum might be infinite (see Section 4.6) or finite: in this last case the supremum might be attained (see Section 4.2) or not (see Section 4.5). Finally, we show that there are classes where $\lambda_{w}$ is constant among its members (see Section 4.7).
4.1. Does every equivalence class have a BRW with pure global survival phase?
The answer is negative. Let us start with a heuristic reasoning before moving on to a more formal proof. Just take $(\mathbb{Z},K)$ where $k_{xy}:=1/2$ if $|x-y|=1$ and 00 otherwise. Consider a modification restricted to an interval $I$. Conditioned to global survival either an infinite number of particles visits $I$ or an infinite number of particles visits $\mathbb{Z}\setminus I$ but in this case, by a Borel-Cantelli argument the progenies of an infinite number of particles visits $I$. So there is no pure global survival phase.
This can be proven rigorously in general as follows. We recall that $(X,K)$ is quasi-transitive if there exists a finite $X_{0}\subset X$ such that for any $x\in X$ there is a bijective map $\gamma_{x}:X\to X$ satisfying $\gamma_{x}^{-1}(x)\in X_{0}$ and $k_{yz}=k_{\gamma_{x}y\,\gamma_{x}z}$ for all $y,z$. For instance, if the rates are translation invariant, then $(X,K)$ is quasi-transitive (actually, it is transitive).
Lemma 4.3.
Let $(X,K)$ be a quasi-transitive, irreducible BRW on an infinite set $X$. Suppose that for all $x\in X$ the set $\{y\in X\colon(x,y)\in E_{\mu}\}$ is finite. If $(X,K)$ has no pure global survival phase, then every BRW in the same class has no pure global survival phase.
Proof.
Here are the main steps of the proofs. Given a subset $A\subseteq X$, we denote by $\lambda_{w}(A)$ and $\lambda_{s}(A)$ the critical parameters of the BRW $(A,K_{A})$ where $K_{A}$ is the restriction of $K$ to $A\times A$ (this is the BRW restricted to $A$). Clearly $\lambda_{w}(A)\geq\lambda_{w}$ and $\lambda_{s}(A)\geq\lambda_{s}$. Moreover, according to [6, Theorem 3.1] (note that the irreducibility of $(A_{n},K_{A_{n}})$ is not needed, see also [23, theorem 5.2]), if $A_{n}\subseteq A_{n+1}$ and $\bigcup_{n}A_{n}=X$ then $\lambda_{s}(A_{n})\downarrow\lambda_{s}$; due to the quasi-transitivity of the BRW, there is $x_{0}\in X$ such that $\{\gamma(x_{0})\colon\gamma\in\mathrm{AUT}(X,K)\}$ is infinite; clearly, for every automorphism $\gamma$ we have $\lambda_{w}(\gamma(A_{n}))=\lambda_{w}(A_{n})$, $\lambda_{s}(\gamma(A_{n}))=\lambda_{s}(A_{n})$. Let us choose $A_{n}$ finite for every $n$; observe that, for every fixed finite $I\subset X$, since $X$ is infinite and the BRW is quasi-transitive, there exists an automorphism $\gamma_{n}$ such that $\gamma_{n}(A_{n})\cap I=\emptyset$. Indeed let $r:=\max\{d(x_{0},y)\colon y\in A_{n}\}$ where $d(x,y):=\min\{n\colon x\stackrel{{\scriptstyle n}}{{\to}}y\}$ (which is not, in general a distance because it might not be symmetric, but it satisfies the triangular inequality). Let the automorphism $\gamma_{n}$ be such that $d(\gamma_{n}(x_{0}),I)>r$, whose existence is guaranteed by the fact that $X$ is infinite and the set $\{y\in X\colon(x,y)\in E_{\mu}\}$ is finite. Thus, $\gamma_{n}(A_{n})\cap I=\emptyset$.
Let $K^{\prime}\in[K]$ such that the matrices are the same outside $I$. Clearly if $\lambda_{s}^{\prime}\leq\lambda_{s}$ or $\lambda_{w}^{\prime}\leq\lambda_{w}$ then $\lambda_{s}^{\prime}=\lambda_{w}^{\prime}$ by Corollary 3.3. Let us suppose that $\lambda_{w}=\lambda_{s}\leq\lambda_{w}^{\prime}\leq\lambda_{s}^{\prime}$. Clearly, $\lambda_{s}(A_{n})\geq\lambda_{s}^{\prime}$ since $\gamma_{n}(A_{n})\cap I=\emptyset$ and $\lambda_{s}(\gamma_{n}(A_{n}))=\lambda_{s}(A_{n})$. Thus
$\lambda_{w}=\lambda_{s}\leq\lambda_{w}^{\prime}\leq\lambda_{s}^{\prime}\leq\lambda_{s}(A_{n})\to\lambda_{s},\qquad\textrm{as }n\to\infty$
therefore $\lambda_{w}=\lambda_{s}=\lambda_{w}^{\prime}=\lambda_{s}^{\prime}$. ∎
4.2. Is there a class without BRWs with pure global survival phase where the supremum of $\lambda_{w}$ is attained?
The answer to this question is affirmative: it is possible to construct a BRW with no pure global survival phase, where all BRWs in its equivalence class has no pure global survival phase, and the supremum value of $\lambda_{w}$ is attained. Consider the class of the BRW $(\mathbb{Z},K)$ where $k_{xy}=1/2$ if $|x-y|=1$ and 00 otherwise. It is easy to prove that $1=\lambda^{K}_{w}=\lambda^{K}_{s}$: the first equality is trivial, while the second one follows, for instance, from [3, Theorem 4.3]. Note that if we consider the same BRW restricted to $\mathbb{N}$, the same equalities $\lambda_{w}(\mathbb{N})=\lambda_{s}(\mathbb{N})=1$ hold: this follows from [3, Theorem 4.3], [23, Theorem 5.2] and from the fact that for a restricted BRW the critical parameters cannot be smaller than those of the original BRW. By Lemma 4.3, there are no BRWs with pure global survival phase in this class.
We prove now that the supremum of $\lambda_{w}$ in the equivalence class of $(\mathbb{Z},K)$ is $1$. Consider $K^{*}\in[K]$ and suppose that the modification is contained in $[-n,+n]$. Thus, by a coupling argument, $\lambda_{w}^{K^{*}}=\lambda_{s}^{K^{*}}\leq\lambda_{w}(\mathbb{N})=1$: indeed $(\mathbb{Z},K^{*})$ dominates its restriction to $[n+1,+\infty)$ which is equivalent to the restriction of $(\mathbb{Z},K)$ to $\mathbb{N}$. To prove that $\lambda_{w}^{K^{*}}$ is not constant in the class, we note that, by adding a loop in 00 with sufficiently large rate $k_{00}$, one can obtain a BRW in this class whose critical values $\lambda_{w}=\lambda_{s}<1$.
4.3. Does every equivalence class have a BRW with no pure global survival phase?
The answer is affirmative. Consider a BRW $(X,K)$ and denote by $\lambda_{w}$ its global critical parameter. Let $o\in X$ be fixed and define $K^{\prime}$ by
$k^{\prime}_{xy}:=k_{xy}+\alpha\delta_{xo}\delta_{yo}.$
Roughly speaking, we modify the just rate from $o$ to itself by adding $\alpha$ to $k_{oo}$. Clearly $\lambda_{s}^{\prime}\leq 1/\alpha$ therefore by taking $\alpha>1/\lambda_{w}$ we have $\lambda^{\prime}_{s}<\lambda_{w}$ thus, according to Corollary 3.3, $\lambda^{\prime}_{w}=\lambda^{\prime}_{s}$.
4.4. If there is a BRW with pure global survival phase in a class, can a BRW with no pure global survival phase have the same (maximum) $\lambda_{w}$?
The answer is affirmative. We rely on Example 4.2. It suffices to note that the BRW on $\mathbb{T}_{d}$ where $k_{xy}=1$ if $x$ and $y$ are neighbours and $k_{xy}=0$ otherwise, has a pure global survival phase, while the modified BRW which adds a loop at the origin $o$, of rate $k_{oo}=\frac{d(d-2)}{d-1}$, has the same global critical parameter $1/d$, but no global survival phase.
4.5. If a class has no BRWs with pure global survival phase and the supremum of $\lambda_{w}$ is finite, is it always attained?
The answer is negative. Consider the class of $(\mathbb{Z},K)$ where $k_{xy}=k_{|x|}/2$ if $|x-y|=1$ and $k_{n}\downarrow\bar{k}$ ad $n\to\infty$. In this case the supremum of $\lambda_{w}$ equals $1/\bar{k}$, indeed take $K^{*}$ as $k^{*}_{xy}:=\min(k_{xy},(\bar{k}+\varepsilon)/2)$ where $\varepsilon>0$; since $\varepsilon$ is a generic positive constant, we have that $\lambda_{w}^{*}\geq 1/(\bar{k}+\varepsilon)$ (since the restriction of a BRW to a subset gives a BRW with larger critical parameters). Then the supremum is at least $1/\bar{k}$. The reversed inequality will be proven in the specific example we are going to construct. The sequence $\{k_{n}\}_{n\in\mathbb{N}}$ will be made of long stretches of finite consecutive constant rates $\bar{k}+1/n$. More specifically, according to [23, Theorem 5.2], for any $n\in\mathbb{N}$ there is $l_{n}\in\mathbb{N}$ such that the BRW on $\mathbb{Z}$ with rates $k_{xy}:=(\bar{k}+1/n)/2$ if $|x-y|=1$ and 00 otherwise (where $\lambda_{s}=\lambda_{w}=1/(\bar{k}+1/n)$), restricted to a set $\{i+1,\ldots,i+l_{n}\}$ has a local critical parameter $\lambda_{s}^{n}$ close to $1/(\bar{k}+1/n)$, say $\lambda_{s}^{n}<1/\bar{k}$. This does not depend on $i$. Define $\{i_{n}\}$ recursively as $i_{n+1}:=i_{n}+l_{n+1}$, where $i_{0}:=0$, and $k_{i}:=\bar{k}+1/(n+1)$ for all $i=i_{n},i_{n}+1,\ldots,i_{n+1}-1$. Consider $K^{*}\in[K]$; clearly there exists $n\in\mathbb{N}$ such that $k^{*}_{i}:=\bar{k}+1/(n+1)$ for all $i=i_{n},i_{n}+1,\ldots,i_{n+1}-1$, therefore $\lambda_{w}^{*}\leq\lambda_{s}^{*}\leq\lambda_{s}^{n}<1/\bar{k}$. This shows that the supremum of $\lambda_{w}$ in the class $[K]$ is $1/\bar{k}$ and it is never attained.
4.6. Is the supremum of $\lambda_{w}$ necessarily finite in a class?
[…]
4.7. Can $\lambda_{w}$ be constant in a class?
The answer is affirmative. Clearly, according to Section 4.3 and Corollary 3.3, the class has no BRWs with pure global survival phase. Take $(\mathbb{N},K)$ where $k_{xy}=k_{|x|}/2$ if $|x-y|=1$ and $k_{n}\uparrow\bar{k}$ ad $n\to\infty$. By spatial approximation ([3, Theorem 4.3], [23, Theorem 5.2]) and domination, $1/\bar{k}\leq\lambda_{w}^{K}\leq\lambda_{s}^{K}\leq 1/\bar{k}$. Since by spatial approximation $\lambda_{s}^{K^{*}}=1/k$ for any modification, then according to Corollary 3.3 then $\lambda_{w}^{K^{*}}=\lambda_{w}^{K}$ (since $\lambda_{s}^{K^{*}}\geq\lambda_{w}^{K}$).